# 1.27in1.27in1in1in

### v1

@@ 4.2. Type II flops on curve structures
${C^{\prime}}^{2}=\pi^{*}{C^{\prime}}^{2}=(C+2E)^{2}=C^{2}+4C.E+4E^{2}=-1+8-4=3.$
By symmetry, the same holds for the anticanonical curve on $Y^{\prime}_{3}$. The squares of the curves in the anticanonical cycle on the special component are given as follows. The two blown up curves correspond to $(-1)$-curves and the other two anticanonical vertices are $(-3)$-curves by the fact that the squares of the glued curves have to add up to zero as soon as one of them is nodal. This last claim is obvious for $Y_{T}$ and is preserved under type I flops, so it holds for our surface $Y^{\prime}$. For simple normal crossing surfaces, this would also follow from the triple point formula, see [FS86, p. 2].
@@
Figure 11 illustrates the type II flop $Y_{P}\dashrightarrow Y^{\prime}$. On each of two components of $Y_{P}$, a double curve is blown down. These components become nonspecial components in the resulting class $T$ surface $Y^{\prime}$. The curves adjacent to the exceptional curves in the curve structure become square zero curves on the corresponding components of $Y^{\prime\nu}$ and are no longer part of the curve structure. The third component of $Y_{P}$, call it $Y_{1}$, becomes the special component $Y_{1}^{\prime}$ of $Y^{\prime}$. Here, the two singular points of the anticanonical cycle on ${\mathfrak{Y}}_{2}$ are blown up. In the curve structure of $Y^{\prime}$, they become the blue anticanonical vertices in the bottom right graph of Figure 11. Lemma 3.9 describes a fibration for which there are uniquely determined smooth rational curves of square zero passing through each blow-up point. They become $(-1)$-curves under the blow-up and become elements in the curve structure of the special component of $Y^{\prime}$. These are the blue shaded curves in the figure.
@@ 5.1. Type I flops on curve structures
[…]
@@
The effect of a type I flop on the curve structure is easy to describe. We flop in a $(-1)$-curve $b$ meeting an anticanonical vertex $d$ (necessarily with multiplicity $1$). This anticanonical vertex is glued to an anticanonical vertex $d^{\prime}$, possibly on a different component. During the flop, the vertex $b$ disappears from the diagram and reappears as a vertex $b^{\prime}$ of square $(-1)$ next to the anticanonical vertex $d^{\prime}$, see Figure 12. For all vertices meeting $b$ before the flop, both their self-intersection and their intersection multiplicity with $d$ increase by one. Conversely, all vertices meeting $d^{\prime}$ before the flop will meet $b^{\prime}$ after the flop, their self-intersection and their intersection multiplicity with $d^{\prime}$ will decrease by one.
@@ 5.2. Curves on combinatorial K3 surfaces
First, we will show that the claim is stable under type I flops. Let $\phi:Y^{\prime}\dashrightarrow Y$ be a type I flop. On the irreducible components, this is either a blow-up, a blow-down, or an isomorphism in which case there is nothing to show. We will show that the claim for $Y$ follows componentwise from the claim for $Y^{\prime}$. Along the blow-down component, say $\phi:Y_{1}^{\prime}\xrightarrow{\ \ }Y_{1}$, this is obvious since the image of a ${\mathbb{Z}}$-basis for $\operatorname{Pic}(Y^{\prime}_{1})$ will constitute a ${\mathbb{Z}}$-basis for $\operatorname{Pic}(Y_{1})$. On the blown up component, say $\phi^{-1}:Y_{2}\xrightarrow{\ \ }Y_{2}^{\prime}$, by the description of Section 5.1 the exceptional curve is added to the curve structure $\Gamma_{Y_{2}}$ and the other vertices of $\Gamma_{Y_{2}}$ correspond to the strict transforms of the vertices in $\Gamma_{Y^{\prime}_{2}}$. Thus, together these vertices still form a ${\mathbb{Z}}$-basis for $\operatorname{Pic}(Y_{2})$.
@@
Let us start with ${\mathfrak{Y}}_{1}$. First take the $(1,3,1,5)$-blow-up of ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ with its toric boundary. Then we successively blow down the $(-1)$-curves in the boundary. Keeping track of the generators, we prove the claim.
Recall that ${\mathfrak{Y}}_{2}$ is obtained from the $(3,3,2)$-blow-up ${\widetilde{{\mathfrak{Y}}}}_{2}$ of ${\mathbb{P}}^{2}$ by contracting the $(-1)$-curve in the toric boundary. As an iterated blow-up, the Picard group of ${\widetilde{{\mathfrak{Y}}}}_{2}$ is generated by the exceptional curves and the strict transform of a line in ${\mathbb{P}}^{2}$. These are depicted in Figure 13, again with the convention that unlabeled vertices represent $(-2)$-curves.
11-1- 111-1- $-1$
Figure 13. Generators of the Picard group for the $(3,3,2)$-blow-up ${\widetilde{{\mathfrak{Y}}}}_{2}$ of ${\mathbb{P}}^{2}$.
The vertex at the fork is the strict transform of a line. The exceptional curve of the blow-down ${\widetilde{{\mathfrak{Y}}}}_{2}\xrightarrow{\ \ }{\mathfrak{Y}}_{2}$ is the upmost black vertex, call it $d$. We have to argue that the strict transforms of the (white) curves not meeting this vertex generate $\operatorname{Pic}({\mathfrak{Y}}_{2})$ over ${\mathbb{Z}}$. Let $v$ denote the unique vertex meeting $d$. Then an intersection number consideration gives $d=-v+w$ with an integral linear combination $w$ of the other white vertices. Thus, in $\operatorname{Pic}({\mathfrak{Y}}_{2})$, the strict transform of $v$ becomes an integral linear combination of the other vertices, which form the curve structure of ${\mathfrak{Y}}_{2}$.
@@ 5.3. The cone of curves
[…]
@@
The statement is verified component by component, so we denote $Y_{1}\subset Y^{\nu}$ a connected component. By [AL11, Theorem 2.1], we know that the cone ${\overline{\operatorname{NE}}}(Y_{1})$ is generated by the closure of the classes of curves of negative square. Note that the assumption $\kappa(-K_{Y_{1}})\geq 0$ is satisfied, as $Y_{1}$ comes with an anticanonical cycle. It thus suffices to show that all curves of negative square show up in the augmented curve structure. Using Lemma 5.4 below, we reduce to the case where $Y_{1}$ is the $n$-blow-up of one of the surfaces ${\mathfrak{Y}}_{i}$. Indeed, if the statement holds for the source of the contraction, it holds for the target by pushforward. Let $C\subset Y_{1}$ be a curve of negative square and let us denote by $D$ the anticanonical cycle. Then either $C\subset D$ or, by our reduction step and [HL22, Proposition 4.5], it is in the curve structure. ∎
@@
[…]
@@
[…]
@@
Proposition 5.3 looks harmless but it is actually quite surprising. Recall that the cone of curves of a rational surface need not be rational polyhedral. In fact, already for the blow-up of ${\mathbb{P}}^{2}$ in $\geq 9$ general points it is not since there are infinitely many $(-1)$-curves by a result of Nagata [Nag61, Theorem 4a]; see also [CLO16, Theorem 2.7] and references therein. The above proposition on the other hand applies to arbitrarily high Picard rank surfaces. The reason is that we are not blowing up in general points but instead in rather special points, the interior special points.
@@
[…]
@@ 6. Flops on Picard
This section is the core of the paper. The maps described here are completely explicit and thus allow one to compute the Mori fan on a computer. We know by Lemma 5.1 that the vertices in the curve structure provide a basis of the Picard group of the normalization of a combinatorial K3 surface. For each flop $\phi:Y\dashrightarrow Y^{\prime}$ of type I or II, we define a map $\phi_{c}:\operatorname{Pic}(Y^{\nu})\xrightarrow{\ \ }\operatorname{Pic}(Y^{\prime\nu})$ between the corresponding Picard groups of the normalizations, see Definitions 6.4 and 6.14. It is referred to as the combinatorial pushforward. The main result of this section is Theorems 6.9 and 6.15, which link the combinatorial pushforward to the pushforward along a flop of the corresponding models of the DNV family, see also Section 6.1. We prove the main theorem separately for type I flops and type II flops.
@@ 6.1. Motivation
(6.1) ${\operatorname{Pic}({\mathcal{Y}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{*}}$$\scriptstyle{\iota^{*}}$${\operatorname{Pic}(Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$?$\scriptstyle{\nu^{*}}$${\operatorname{Pic}(Y^{\nu})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{c}}$${\operatorname{Pic}({\mathcal{Y}}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota^{\prime*}}$${\operatorname{Pic}(Y^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu^{*}}$${\operatorname{Pic}(Y^{\prime\nu})}$
where $\iota$, $\iota^{\prime}$ denote the inclusions of the central fibers. The statement of Theorem 6.9 resp. Theorem 6.15 is that the outer diagram is commutative in the degree $22d=$ 2 case. In particular, the combinatorial pushforward $\phi_{c}$ restricts to $\iota^{\prime*}\circ\Phi_{*}\circ\left(\iota^{*}\right)^{-1}$ on $\operatorname{Pic}(Y)$. This is remarkable for at least two reasons.
@@ 6.2. Type I Flops on Picard
Setup 6.1.
Let $\phi:Y\dashrightarrow Y^{\prime}$ be a type I flop between combinatorial K3 surfaces. We denote by ${\mathcal{B}}_{\operatorname{std}},{\mathcal{B}}_{\operatorname{std}}^{\prime}$ the standard bases of $Y,Y^{\prime}$. Then there are vertices $v_{\operatorname{flop}}\in{\mathcal{B}}_{\operatorname{std}}$, $v_{\operatorname{flop}}^{\prime}\in{\mathcal{B}}_{\operatorname{std}}^{\prime}$ such that (the curve corresponding to) $v_{\operatorname{flop}}$ is contracted under $\phi$ and (the curve corresponding to) $v_{\operatorname{flop}}^{\prime}$ is contracted under $\phi^{-1}$. We will denote $(\cdot)^{\prime}:{\mathcal{B}}_{\operatorname{std}}\xrightarrow{\ \ }{\mathcal{B}}_{\operatorname{std}}^{\prime}$ be the bijection sending $v_{\operatorname{flop}}$ to $v_{\operatorname{flop}}^{\prime}$ and all other vertices to their strict transforms. We denote the intersection pairings on $Y,Y^{\prime}$ by $(\cdot,\cdot),(\cdot,\cdot)^{\prime}$.
@@
(6.2) ${\mathcal{B}}_{\operatorname{std}}={\mathcal{B}}_{c}\cup{\mathcal{B}}_{\operatorname{old}}\cup{\mathcal{B}}_{\operatorname{new}}\cup\{v_{\operatorname{flop}}\}\quad\textrm{ and }\quad{\mathcal{B}}_{\operatorname{std}}^{\prime}={\mathcal{B}}_{c}^{\prime}\cup{\mathcal{B}}^{\prime}_{\operatorname{old}}\cup{\mathcal{B}}^{\prime}_{\operatorname{new}}\cup\{v^{\prime}_{\operatorname{flop}}\}$
@@
[…]
@@
$\left(\phi_{c}(v_{\operatorname{flop}}),\phi_{c}(v)\right)^{\prime}$ $=(-v_{\operatorname{flop}}^{\prime},v^{\prime}+v_{\operatorname{flop}}^{\prime})^{\prime}=1-(v_{\operatorname{flop}}^{\prime},v^{\prime})^{\prime}=\begin{cases}0&\textrm{ if }v^{\prime}\in{\mathcal{B}}^{\prime}_{\operatorname{new}}\\ 1&\textrm{ if }v^{\prime}\in{\mathcal{B}}^{\prime}_{\operatorname{old}}\end{cases}$
@@
We will thus identify ${\widetilde{{\mathcal{Y}}}}:={\operatorname{Bl}}_{B}{\mathcal{Y}}={\operatorname{Bl}}_{B^{\prime}}{\mathcal{Y}}^{\prime}$. Then the exceptional loci of $p$ and $q$ coincide and are given by an irreducible divisor $E{\ \cong\ }{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$. Under this identification, $p$ and $q$ are identified with the projections to the factors. Moreover, $B^{\prime}$ is the image under $q$ of a fiber of $p|_{E}$. Then the linear map $\Phi_{*}$ is given by $q_{*}p^{*}$.∎
Before we come to the main result of this section, we will treat a special case, which can be very instructive.
@@
We thus have to find an effective divisor $D\subset{\mathcal{Y}}_{P}$ whose restriction $D_{0}$ to the central fiber is $A$ resp. $B+C$ and show that in this case $D^{\prime}:=\Phi_{*}D$ restricts to $A^{\prime}+B^{\prime}$ resp. $C^{\prime}$ on the central fiber of ${\mathcal{Y}}^{\prime}$. In both cases, $D$ is unique: the classes in $\operatorname{Pic}({\mathcal{Y}}_{P})$ extending $a$ resp. $b+c$ both have square $(-2)$ and thus are either effective or anti-effective. The latter alternative is excluded, since an effective class must specialize to an effective class. Let us denote by $D$ an effective divisor on ${\mathcal{Y}}_{P}$ representing these classes, one sees that it must be irreducible. Indeed, if there were several irreducible components, all would restrict to $(-2)$ curves on the K3 surface $({\mathcal{Y}}_{P})_{\eta}$ so they cannot restrict to $A$ or $B+C$ on the central fiber.
@@
While the geometric description of $\Phi_{*}$ was not too difficult in the special case of Example 6.7, it can get rather involved in general, mainly due to the various possibilities for the curve structure in a neighborhood of the flopped curve. Instead, we will undertake a different approach based on intersection products and the projection formula. For this, let ${\widetilde{K}}\subset{\widetilde{{\mathcal{Y}}}}:={\operatorname{Bl}}_{B}{\mathcal{Y}}_{P}$ be an integral curve and denote $K:=p_{*}{\widetilde{K}}$ and $K^{\prime}:=q_{*}{\widetilde{K}}$. By the projection formula, for every divisor $D$ on ${\mathcal{Y}}_{P}$ we have
@@
(6.4) $D_{0}$ $\sim\sum_{v\in{\mathcal{B}}_{\operatorname{old}}}a_{v}A_{v}+\sum_{v\in{\mathcal{B}}_{\operatorname{new}}}b_{v}B_{v}+\sum_{v\in{\mathcal{B}}_{c}}c_{v}C_{v}+\lambda B\quad\textrm{ resp. }$.
@@
00 0 $=D.p_{*}F^{\prime}=p^{*}D.F^{\prime}={\widetilde{D}}.F^{\prime}$
@@
00 0 $=D^{\prime}.q_{*}F=q^{*}D^{\prime}.F={\widetilde{D}}.F+\mu^{\prime}E.F$

### v2

@@ 4.2. Type II flops on curve structures
${C^{\prime}}^{2}=(\pi^{*}{C^{\prime}})^{2}=(C+2E)^{2}=C^{2}+4C.E+4E^{2}=-1+8-4=3.$
By symmetry, the same holds for the anticanonical curve on $Y^{\prime}_{3}$. The squares of the curves in the anticanonical cycle on the special component are given as follows. The two blown up curves correspond to (1)(-1)-curves and the other two anticanonical vertices are (3)(-3)-curves by the fact that the squares of the glued curves have to add up to zero as soon as one of them is nodal. This last claim is obvious for $Y_{T}$ and is preserved under type I flops, so it holds for our surface $Y^{\prime}$. For simple normal crossing surfaces, this would also follow from the triple point formula, see [FS86, p. 2].
@@
Figure 11 illustrates the type II flop $Y_{P}\dashrightarrow Y^{\prime}$. On each of two components of $Y_{P}$, a double curve is blown down. These components become nonspecial components in the resulting class $T$ surface $Y^{\prime}$. The curves adjacent to the exceptional curves in the curve structure become square zero curves on the corresponding components of $Y^{\prime\nu}$ and are no longer part of the curve structure. The third component of $Y_{P}$, call it $Y_{1}$, becomes the special component $Y_{1}^{\prime}$ of $Y^{\prime}$. Here, the two singular points of the anticanonical cycle on ${\mathfrak{Y}}_{2}$ are blown up. In the curve structure of $Y^{\prime}$, they become the blue anticanonical vertices in the bottom right graph of Figure 11. Lemma 3.9 describes a fibration for which there are uniquely determined smooth rational curves of square zero passing through each blow-up point. They become (1)(-1)-curves under the blow-up and become elements in the curve structure of the special component of $Y^{\prime}$. These are the blue shaded curves in the figure.
@@ 5.1. Type I flops on curve structures
$\ldots\ldots$$\ldots\ldots1-11-12-2$type I flop
@@
The effect of a type I flop on the curve structure is easy to describe. We flop in a (1)(-1)-curve $b$ meeting an anticanonical vertex $d$ (necessarily with multiplicity $1$). This anticanonical vertex is glued to an anticanonical vertex $d^{\prime}$, possibly on a different component. During the flop, the vertex $b$ disappears from the diagram and reappears as a vertex $b^{\prime}$ of square (1)(-1) next to the anticanonical vertex $d^{\prime}$, see Figure 12. For all vertices meeting $b$ before the flop, both their self-intersection and their intersection multiplicity with $d$ increase by one. Conversely, all vertices meeting $d^{\prime}$ before the flop will meet $b^{\prime}$ after the flop, their self-intersection and their intersection multiplicity with $d^{\prime}$ will decrease by one.
@@ 5.2. Curves on combinatorial K3 surfaces
For the proof and for the remainder of this section, recall the notion of an $n$-blow-up from Definition 2.16.
@@
First, we will show that the claim is stable under type I flops. Let $\phi:Y^{\prime}\dashrightarrow Y$ be a type I flop. On the irreducible components, this is either a blow-up, a blow-down, or an isomorphism in which case there is nothing to show. We will show that the claim for $Y$ follows componentwise from the claim for $Y^{\prime}$. Along the blown-down component, say $\phi:Y_{1}^{\prime}\xrightarrow{\ \ }Y_{1}$, this is obvious since the image of a $\mathbb{Z}$}-basis for $\operatorname{Pic}(Y^{\prime}_{1})$ will constitute a $\mathbb{Z}$}-basis for $\operatorname{Pic}(Y_{1})$. On the blown up component, say $\phi^{-1}:Y_{2}\xrightarrow{\ \ }Y_{2}^{\prime}$, by the description of Section 5.1 the exceptional curve is added to the curve structure $\Gamma_{Y_{2}}$ and the other vertices of $\Gamma_{Y_{2}}$ correspond to the strict transforms of the vertices in $\Gamma_{Y^{\prime}_{2}}$. Thus, together these vertices still form a $\mathbb{Z}$}-basis for $\operatorname{Pic}(Y_{2})$.
@@
Let us start with ${\mathfrak{Y}}_{1}$. By Example 2.17, it is the $8$-blow-up of ${\mathbb{P}}^{2}$ in an inflection point of a nodal cubic. Thus, the Picard group of ${\mathfrak{Y}}_{1}$ is generated by the exceptional curves and the strict transform of a line in ${\mathbb{P}}^{2}$. These are precisely the curves in the curve structure.
Recall from Example 2.18 that ${\mathfrak{Y}}_{2}$ is obtained as a $(3,3)$-blow-up of the Hirzebruch surface ${\mathbb{F}}_{2}$. Thus, the Picard group of ${\mathfrak{Y}}_{2}$ is generated by the exceptional curves and the strict transform of a fiber and the (2)(-2)-curve on ${\mathbb{F}}_{2}$. Again, these are precisely the curves in the curve structure.
@@ 5.3. The cone of curves
Let $Y\in{\operatorname{CK3}}$ and let $Y^{\nu}$ be its normalization. Then the cone ${\overline{\operatorname{NE}}}(Y^{\nu})$ is generated by all curves in the augmented curve structure, i.e. the curve structure plus the anticanonical vertices. In particular, ${\overline{\operatorname{NE}}}(Y^{\nu})$ is rational polyhedral.
@@
The statement is verified component by component, so we denote $Y_{1}\subset Y^{\nu}$ a connected component. Let us first assume that $Y_{1}$ has Picard rank $\rho(Y_{1})\geq 3$. In this case, by [AL11, Theorem 2.1], we know that the cone ${\overline{\operatorname{NE}}}(Y_{1})$ is generated by the closure of the classes of curves of negative square. Note that the assumption $\kappa(-K_{Y_{1}})\geq 0$ is satisfied, as $Y_{1}$ comes with an anticanonical cycle. It thus suffices to show that all curves of negative square show up in the augmented curve structure. This is true by Corollary 5.5 below.
We turn to the case where $Y_{1}$ has Picard rank at most $2$. Such a surface is obtained as a birational contraction of a surface of the same class having Picard rank at least $3$. Let $\pi:Y_{1}^{\prime}\xrightarrow{\ \ }Y_{1}$ be such a contraction. Hence, the claim holds for $Y_{1}^{\prime}$ by what was proven so far. By definition of the augmented curve structure (see Definition 2.23), we obtain the one of $Y_{1}$ from the one of $Y_{1}^{\prime}$ by pushforward along $\pi$. Hence, given a curve $C\subset Y_{1}$, we can write the class of $\pi^{-1}(C)\subset Y_{1}^{\prime}$ as a linear combination with nonnegative coefficients of the curves in the curve structure $\Gamma_{Y^{\prime}_{1}}^{a}$. Then the class of $C$ is equal to the pushforward of this linear combination, which, by what we have just said and the definition of a curve structure, is a linear combination with nonnegative coefficients of the curves in the curve structure $\Gamma_{Y_{1}}^{a}$. The claim follows. ∎
@@
Let $Y_{1}$ be a component of the normalization of a surface in $\operatorname{CK3}$} and denote by $D_{1},\ldots,D_{m}$ the components of its anticanonical cycle. Put $n_{i}:=\max(0,D_{i}^{2}+1)$ and $n:=(n_{1},\ldots,n_{m})$. Further, put ${\mathfrak{n}}_{i}:=\max(0,-1-D_{i}^{2})$ and ${\mathfrak{n}}:=({\mathfrak{n}}_{1},\ldots,{\mathfrak{n}}_{m})$. Then there are morphisms
@@
Let $Y_{1}$ be a component of a surface in $\operatorname{CK3}$}. Then all square-negative curves on $Y_{1}^{\nu}$ appear in the curve structure.
Proof.
Using Lemma 5.4, we reduce to the case where $Y_{1}$ is the $n$-blow-up of some ${\mathfrak{Y}}_{i}$. Indeed, if the statement holds on a blow-up, it holds on the surface itself by pushforward. Let $C\subset Y_{1}$ be a curve of negative square and let us denote by $D$ the anticanonical cycle. Then either $C\subset D$ or, by our reduction step and [HL22, Proposition 4.5], it is in the curve structure. ∎
@@
Proposition 5.3 looks harmless, but it is actually quite surprising. Recall that the cone of curves of a rational surface need not be rational polyhedral. In fact, already for the blow-up of ${\mathbb{P}}^{2}$ in $\geq 9$ general points it is not since there are infinitely many (1)(-1)-curves by a result of Nagata [Nag61, Theorem 4a]; see also [CLO16, Theorem 2.7] and references therein. The above proposition on the other hand applies to arbitrarily high Picard rank surfaces. The reason is that we are not blowing up in general points but instead in rather special points, the interior special points.
@@
Let $Y\in{\operatorname{CK3}}$. By (2.1), we know how to describe the Picard group $\operatorname{Pic}(Y)$ inside $\operatorname{Pic}(Y^{\nu})$. Proposition 5.3 allows us to compute the nef cone ${\operatorname{Nef}}(Y^{\nu})$ as the dual of the cone of curves. One can algorithmically compute the ray generators of a rational polyhedral cone, given generators for its dual cone. This is well known and implemented, for example, in SageMath [Sage24]. Finally, this allows to compute ${\operatorname{Nef}}(Y)$ as
@@ 6. Flops on Picard
This section is the core of the paper. The maps described here are completely explicit and thus allow one to compute the Mori fan on a computer. We know by Lemma 5.1 that the vertices in the curve structure provide a basis of the Picard group of the normalization of a combinatorial K3 surface. For each flop $\phi:Y\dashrightarrow Y^{\prime}$ of type I or II, we define a map $\phi_{c}:\operatorname{Pic}(Y^{\nu})\xrightarrow{\ \ }\operatorname{Pic}(Y^{\prime\nu})$ between the corresponding Picard groups of the normalizations, see Definitions 6.4 and 6.14. It is referred to as the combinatorial pushforward. The main results of this section are Theorems 6.9 and 6.15, which link the combinatorial pushforward to the pushforward along a flop of the corresponding models of the DNV family, see also Section 6.1. We prove the main theorem separately for type I flops and type II flops.
@@ 6.1. Motivation
(6.1) ${\operatorname{Pic}({\mathcal{Y}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{*}}$$\scriptstyle{\iota^{*}}$${\operatorname{Pic}(Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu^{*}}$${\operatorname{Pic}(Y^{\nu})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{c}}$${\operatorname{Pic}({\mathcal{Y}}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota^{\prime*}}$${\operatorname{Pic}(Y^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu^{*}}$${\operatorname{Pic}(Y^{\prime\nu})}$
where $\iota$, $\iota^{\prime}$ denote the inclusions of the central fibers. The statement of Theorem 6.9 resp. Theorem 6.15 is that the outer diagram is commutative in the degree $2d=2$ case. In particular, the combinatorial pushforward $\phi_{c}$ restricts to $\iota^{\prime*}\circ\Phi_{*}\circ\left(\iota^{*}\right)^{-1}$ on $\operatorname{Pic}(Y)$. This is remarkable for at least two reasons.
@@ 6.2. Type I Flops on Picard
Recall from Lemma 5.1 that for $Y\in{\operatorname{CK3}}$, the Picard group $\operatorname{Pic}(Y^{\nu})$ has a canonical $\mathbb{Z}$}-basis given by the vertices of the associated curve structure. We refer to this basis as the standard basis of $Y$.
Setup 6.1.
@@
[…]
@@
(6.3) $\phi_{c}(v):=\begin{cases}-v_{\operatorname{flop}}^{\prime}&\textrm{ if }v=v_{\operatorname{flop}}\\[2.0pt] v^{\prime}+v^{\prime}_{\operatorname{flop}}&\textrm{ if }v\in{\mathcal{B}}_{\operatorname{new}}\cup{\mathcal{B}}_{\operatorname{old}}\\[2.0pt] v^{\prime}&\textrm{ otherwise.}\end{cases}$
@@
[…]
@@
${{\operatorname{Bl}}_{B}{\mathcal{Y}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{{\ \cong\ }}$$\scriptstyle{p}$${{\operatorname{Bl}}_{B^{\prime}}{\mathcal{Y}}^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q}$${{\mathcal{Y}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${{\mathcal{Y}}^{\prime}}$
@@
Before we come to the main result of this section, we will treat a special case, which is very instructive.
@@
We thus have to find an effective divisor $D\subset{\mathcal{Y}}_{P}$ whose restriction $D_{0}$ to the central fiber is $A$ resp. $B+C$ and show that in this case $D^{\prime}:=\Phi_{*}D$ restricts to $A^{\prime}+B^{\prime}$ resp. $C^{\prime}$ on the central fiber of $\mathcal{Y}$$^{\prime}$. In both cases, $D$ is unique: the classes in $\operatorname{Pic}({\mathcal{Y}}_{P})$ extending $a$ resp. $b+c$ both have square (2)(-2) and thus are either effective or anti-effective. The latter alternative is excluded, since an effective class must specialize to an effective class. Let us denote by $D$ an effective divisor on ${\mathcal{Y}}_{P}$ representing these classes, one sees that it must be irreducible. Indeed, if there were several irreducible components, all would restrict to (2)(-2) curves on the K3 surface $({\mathcal{Y}}_{P})_{\eta}$ so they cannot restrict to $A$ or $B+C$ on the central fiber.
@@
While the geometric description of $\Phi_{*}$ is not too difficult in the special case of Example 6.7, it can get rather involved in general, mainly due to the various possibilities for the curve structure in a neighborhood of the flopped curve. Instead, we will undertake a different approach based on intersection products and the projection formula. For this, let ${\widetilde{K}}\subset{\widetilde{{\mathcal{Y}}}}:={\operatorname{Bl}}_{B}{\mathcal{Y}}_{P}$ be an integral curve and denote $K:=p_{*}{\widetilde{K}}$ and $K^{\prime}:=q_{*}{\widetilde{K}}$. By the projection formula, for every divisor $D$ on ${\mathcal{Y}}_{P}$ we have
@@
(6.4) $D_{0}$ $\sim\sum_{v\in{\mathcal{B}}_{\operatorname{old}}}a_{v}A_{v}+\sum_{v\in{\mathcal{B}}_{\operatorname{new}}}b_{v}B_{v}+\sum_{v\in{\mathcal{B}}_{c}}c_{v}C_{v}+\lambda B\quad\textrm{ resp. }$
@@
$0$ $=D.p_{*}F^{\prime}=p^{*}D.F^{\prime}={\widetilde{D}}.F^{\prime}$
@@
$0$ $=D^{\prime}.q_{*}F=q^{*}D^{\prime}.F={\widetilde{D}}.F+\mu^{\prime}E.F$